# Low-Rank Graph Contrastive Learning for Node Classification

### v1

@@ 5.1 Experimental Settings
Rank Noise Level
@@
Tuning $r_{0},\tau$ by Cross-Validation. We tune the rank $r_{0}$ and the weight for the truncated nuclear loss $\tau$ by standard cross-validation on each dataset. Let $r_{0}=\lceil\gamma\min\left\{N,d\right\}\rceil$ where $\gamma$ is the rank ratio. We select the values of $\gamma$ and $\tau$ by performing $5$-fold cross-validation on $20\%$ of the training data in each dataset. The value of $\gamma$ is selected from $\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9\}$. The value of $\tau$ is selected from $\{0.05,0.1,0.15,0.2,0.25,0.3,0.35,0.4,0.45,0.5\}$. The selected values on each dataset are shown in Table 3.
Table 3: Selected rank ratio $\gamma$ and truncated nuclear loss’s weight $\lambda$ for each dataset.
Hyper-parameters Cora Citeseer PubMed Coauthor CS ogbn-arxiv Wiki-CS Amazon-Computers Amazon-Photos
$\tau$ 0.10 0.10 0.10 0.20 0.10 0.25 0.20 0.20
$\gamma$ 0.2 0.2 0.3 0.3 0.4 0.2 0.2 0.3
Table 4: The statistics of the datasets.
Dataset Nodes Edges Features Classes
Cora 2,708 5,429 1,433 7
CiteSeer 3,327 4,732 3,703 6
PubMed 19,717 44,338 500 3
Coauthor CS 18,333 81,894 6,805 15
ogbn-arxiv 169,343 1,166,243 128 40
Wiki-CS 11,701 215,863 300 10
Amazon-Computers 13,752 245,861 767 10
Amazon-Photos 7,650 119,081 745 8
5.2 Datasets
We evaluate our method on eight public benchmarks that are widely used for node representation learning, namely Cora, Citeseer, PubMed [Sen et al., 2008], Coauthor CS, ogbn-arxiv [Hu et al., 2020], Wiki-CS [Mernyei and Cangea, 2020], Amazon-Computers, and Amazon-Photos [Shchur et al., 2018]. Cora, Citeseer, and PubMed are the three most widely used citation networks. Coauthor CS is a co-authorship graph. The ogbn-arxiv is a directed citation graph. Wiki-CS is a hyperlink networks of computer science articles. Amazon-Computers and Amazon-Photos are co-purchase networks of products selling on Amazon.com. We summarize the statistics of all the datasets in Table 4. For all our experiments, we follow the default separation [Shchur et al., 2018, Mernyei and Cangea, 2020, Hu et al., 2020] of training, validation, and test sets on each benchmark.
5.3 Node Classification
[…]
Experimental Results.
We first compare LR-GCL against competing methods for semi-supervised or transductive node classification on input with two types of label noise. To show the robustness of LR-GCL against label noise, we perform the experiments on graphs injected with different levels of label noise ranging from $40\%$ to $80\%$ with a step of $20\%$. We follow the widely used semi-supervised setting [Kipf and Welling, 2017] for node classification. In LR-GCL, we train a transductive classifier for node classification. Previous GCL methods, including MERIT, SUGRL, and SFA, train a linear layer for inductive classification on top of the node representations learned by contrastive learning without using test data in training. Because LR-GCL is a transductive classifier, for fair comparisons, we also train the compared GCL baselines with the same transductive classifier as that for LR-GCL and a two-layer GCN transductive classifier. The results with different types of classifiers are deferred in Section C.2 of the supplementary. For all the baselines in our experiments which perform inductive classification when predicting tbe labels, we report their best results among using their original inductive classifier and two types of transductive classifiers: the same transductive classifier as that for LR-GCL and a two-layer GCN transductive classifier.
Results on Cora, Citeseer, PubMed, Coauthor-CS, and ogbn-arxiv are shown in Table 1, where we report the means of the accuracy of $10$ runs and the standard deviation. Results on Wiki-CS, Amazon-Computers, and Amazon-Photos are deferred in Section C.1 of the supplementary. It is observed from the results that LR-GCL outperforms all the baselines. By selecting confident nodes and computing robust prototypes using BEC, LR-GCL outperforms all the baselines by an even larger margin with a larger label noise level. In addition, we compare LR-GCL with baselines for noisy input with attribute noise levels ranging from $40\%$ to $80\%$ with a step of $20\%$. Results on Cora, Citeseer, and Coauthor CS are shown in Table 5 in the supplementary, where we report the means of the accuracy of $10$ runs and the standard deviation. The results clearly show that LR-GCL is more robust to attribute noise compared to all the baselines for different noise levels.
Additional Results and Ablation Studies We compare the training time of LR-GCL with competing baselines in Table 7 of the supplementary. We also perform ablation study on the value of rank $r_{0}$ in the truncated nuclear norm ${\left\|\mathbf{H}\right\|}_{r_{0}}$ in the loss function (2) of LR-GCL. It is observed from Table 2 that the performance of our LR-GCL is consistently close to the best performance among all the choices of the rank when $r_{0}$ is between $0.1\min\left\{N,d\right\}$ and $0.3\min\left\{N,d\right\}$.
6 Conclusions
In this paper, we propose a novel GCL encoder termed Low-Rank Graph Contrastive Learning (LR-GCL). LR-GCL is a robust GCL encoder which produces low-rank features inspired by the low frequency property of universal graph datasets and the sharp generalization bound for transductive learning. LR-GCL is trained with prototypical GCL with the truncated nuclear norm as the regularization term. We evaluate the performance of LR-GCL with comparison to competing baselines on semi-supervised or transductive node classification, where graph data are corrupted with noise in either the labels for the node attributes. Extensive experimental results demonstrate that LR-GCL generates more robust node representations with better performance than the current state-of-the-art node representation learning methods.
Appendix A Theoretical Results
First, recall that the optimization problem of the low-rank transductive classification is
$\min\limits_{\mathbf{W}}\frac{1}{m}\sum_{v_{i}\in\mathcal{V_{L}}}\textup{KL}\left(\mathbf{y}_{i},\left[\textup{softmax}\left(\mathbf{H}\mathbf{W}\right)\right]_{i}\right).$
We then present the proof of Theorem 4.1.
Proof of Theorem 4.1.
It can be verified that at the $t$-th iteration of gradient descent for $t\geq 1$, we have
$\mathbf{W}^{(t)}=\mathbf{W}^{(t-1)}-\eta\mathbf{H}^{\top}\left(\mathbf{H}\mathbf{W}^{(t-1)}-\mathbf{Y}\right).$ (5)
It follows by (5) that
$\mathbf{H}\mathbf{W}^{(t)}=\mathbf{H}\mathbf{W}^{(t-1)}-\eta\mathbf{K}\left(\mathbf{H}\mathbf{W}^{(t-1)}-\mathbf{Y}\right),$ (6)
where $\mathbf{K}=\mathbf{H}\mathbf{H}^{\top}$. With $\mathbf{F}(\mathbf{W},t)=\mathbf{H}\mathbf{W}^{(t)}$, it follows by (6) that
$\mathbf{F}(\mathbf{W},t)-\mathbf{Y}=\left(\mathbf{I}_{N}-\eta\mathbf{K}\right)\left(\mathbf{F}(\mathbf{W},r,t-1)-\mathbf{Y}\right).$
It follows from the above equality that
$\mathbf{F}(\mathbf{W},t)-\mathbf{Y}=\left(\mathbf{I}_{N}-\eta\mathbf{K}\right)^{t}\left(\mathbf{F}(\mathbf{W},0)-\mathbf{Y}\right)=-\left(\mathbf{I}_{N}-\eta\mathbf{K}\right)^{t}\mathbf{Y},$
so that
${\left\|\mathbf{F}(\mathbf{W},t)-\mathbf{Y}\right\|}_{\textup{F}}\leq{\left\|\left(\mathbf{I}_{N}-\eta\mathbf{K}\right)^{t}\mathbf{Y}\right\|}_{\textup{F}}.$ (7)
As a result of (7), we have
${\left\|\left[\mathbf{F}(\mathbf{W},t)-\mathbf{Y}\right]_{{\cal L}}\right\|}_{\textup{F}}\leq{\left\|\left(\mathbf{I}_{N}-\eta\mathbf{K}\right)^{t}\mathbf{Y}\right\|}_{\textup{F}}.$ (8)
We apply [Yang, 2023, Corollary 3.7] to obtain the following bound for the test loss $\frac{1}{u}{\left\|\left[\mathbf{F}(\mathbf{W},t)-\mathbf{Y}\right]_{{\cal U}}\right\|}_{\textup{F}}^{2}$:
$\frac{1}{u}{\left\|\left[\mathbf{F}(\mathbf{W},t)-\mathbf{Y}\right]_{{\cal U}}\right\|}_{\textup{F}}^{2}$ $\leq\frac{c_{0}}{m}{\left\|\left[\mathbf{F}(\mathbf{W},t)-\mathbf{Y}\right]_{{\cal L}}\right\|}_{\textup{F}}^{2}+c_{1}\min_{0\leq Q\leq n}r(u,m,Q)+\frac{c_{2}x}{u},$ (9)
with
$r(u,m,Q)\coloneqq Q\left(\frac{1}{u}+\frac{1}{m}\right)+\left(\sqrt{\frac{\sum\limits_{q=Q+1}^{n}\widehat{\lambda}_{q}}{u}}+\sqrt{\frac{\sum\limits_{q=Q+1}^{n}\widehat{\lambda}_{q}}{m}}\right).$
where $c_{0},c_{1},c_{2}$ are positive numbers depending on $\mathbf{U}$, $\left\{\widehat{\lambda}_{i}\right\}_{i=1}^{r}$, and $\tau_{0}$ with $\tau_{0}^{2}=\max_{i\in[N]}\mathbf{K}_{ii}$.
It follows by (7) and (9) with $Q=r_{0}$ that
$\frac{1}{u}{\left\|\left[\mathbf{F}(\mathbf{W},t)-\mathbf{Y}\right]_{{\cal U}}\right\|}_{\textup{F}}^{2}$
$\leq\frac{c_{0}}{m}{\left\|\left[\mathbf{F}(\mathbf{W},t)-\mathbf{Y}\right]_{{\cal L}}\right\|}_{\textup{F}}^{2}+c_{1}r_{0}\left(\frac{1}{u}+\frac{1}{m}\right)+\left(\sqrt{\frac{\sum\limits_{q=r_{0}+1}^{n}\widehat{\lambda}_{q}}{u}}+\sqrt{\frac{\sum\limits_{q=r_{0}+1}^{n}\widehat{\lambda}_{q}}{m}}\right)+\frac{c_{2}x}{u},$ (10)
which completes the proof.
Appendix B Implementation Details
B.1 Compared Methods
To demonstrate the power of LR-GCL in learning robust node representation, we compare LR-GCL with two robust contrastive learning baselines, Jo-SRC and Sel-CL, which select clean samples for image data. Since their sample selection methods are general and not limited to the image domain, we adopt these two baselines in our experiments. Next, we introduce the implementation details of applying Jo-SRC and Sel-CL on graph data.
Jo-SRC [Yao et al., 2021]: Jo-SRC is a robust contrastive learning method proposed for image classification. It selects clean samples for training by adopting the Jensen-Shannon divergence to measure the likelihood of each sample being clean. Because this method is a general selection strategy on the representation space, it is adapted to selecting clean samples in the representation space of nodes in GCL. It also introduces a consistency regularization term to the contrastive loss to improve the robustness. To get a competitive robust GCL baseline, we apply the sample selection strategy and the consistency regularization proposed by Jo-SRC to state-of-the-art GCL methods MVGRL, MERIT, and SUGRL. We add the regularization term in Jo-SRC to the graph contrastive loss. The GCL encoders are trained only on the clean samples selected by Jo-SRC. We only report the best results for comparison, which are achieved by applying Jo-SRC to MERIT.
Sel-CL [Li et al., 2022]: Sel-CL is a supervised contrastive learning proposed to learn robust pre-trained representations for image classification. It proposes to select confident contrastive pairs in the contrastive learning frameworks. Sel-CL first selects confident examples by measuring the agreement between learned representations and labels generated by label propagation with cross-entropy loss. Next, Sel-CL selects contrastive pairs from those with selected confident examples in them. This method is also a general sample selection strategy on a learned representation space. So we can adapt Sel-CL to the node representation space to select confident pairs for GCL. In this process, they only select contrastive pairs whose representation similarity is higher than a dynamic threshold. In our experiments, we also adopt the confident contrastive pair selection strategy to the state-of-the-art GCL methods MVGRL, MERIT, and SUGRL. With the same GCL framework, GCL encoders are trained only on the confident pairs selected by Sel-CL. We only report the best results for comparison, which are achieved by applying Sel-CL to MERIT.
Appendix C Additional Experimental Results
C.1 Node Classification on Wiki-CS, Amazon-Computers, and Amazon-Photos
The results for node classification with symmetric label noise, asymmetric label noise, and attribute noise on Wiki-CS, Amazon-Computers, and Amazon-Photos are shown in Table 5. It is observed that LR-GCL also outperforms all the baselines for node classification with both label noise and attribute noise on these three benchmark datasets.
@@ C.1 Node Classification on Wiki-CS, Amazon-Computers, and Amazon-Photos
Table 5: Performance comparison for node classification on Wiki-CS, Amazon-Computers, and Amazon-Photos with asymmetric label noise, symmetric label noise, and attribute noise.
Noise Level
@@
LR-GCL 0.757$\pm$0.010 0.520$\pm$0.013 0.581$\pm$0.013 0.570$\pm$0.007 0.410$\pm$0.014 0.455$\pm$0.014 0.406$\pm$0.012 0.369$\pm$0.012 0.335$\pm$0.014 0.318$\pm$0.010
C.2 Node Classification Results for GCL Methods with different types of classifiers
Existing GCL methods, such as MERIT, SUGRL, and SFA, first train a graph encoder with graph contrastive learning objectives such as InfoNCE [Jin et al., 2021]. After obtaining the node representation learned by contrastive learning, a linear layer for classification is trained in the supervised setting. In contrast, LR-GCL adopts a transductive classifier on top of the node representation obtained by contrastive learning. For fair comparisons with previous GCL methods, we also train the compared GCL baselines with the same transductive classifier as in LR-GCL and a two-layer transductive GCN classifier. The results with different types of classifiers are deferred in Section C.2 of the supplementary.
@@ C.2 Node Classification Results for GCL Methods with different types of classifiers
Table 6: Performance comparison for node classification by inductive linear classifier, transductive two-layer GCN classifier, and transductive classifier used in LR-GCL. The comparisons are performed on Cora with asymmetric label noise, symmetric label noise, and attribute noise.
Noise Level

### v2

@@ 5.5 Ablation Study on the Rank in the Truncated Nuclear Norm
Rank Noise Type
@@
5.6 Additional Study in the Kernel Complexity
In this section, we evaluate the KC of the gram matrix derived from node representations learned by GCL-LRR and other competing GCL methods across various datasets, under symmetric label noise at a corruption level of $40$, as discussed in Theorem 4.1. The corresponding results are reported in Table 7. As observed, the Gram matrices produced by GCL-LRR exhibit substantially lower complexity compared to those of baseline methods. This indicates that transductive classifiers trained on GCL-LRR representations are likely to achieve lower generalization errors on the unlabeled nodes.
Table 7: Comparisons in the kernel complexity defined in Theorem 4.1. The evaluation is performed on the semi-supervised node classification task with $40\%$ of symmetric label noise.
Datasets MERIT SFA Jo-SRC GCN GFSA HONGAT GCL-LRR GCL-LR-Attention
Cora KC 0.37 0.42 0.48 0.44 0.35 0.40 0.20 0.18
$r_{0}$ 1420 1478 1665 1511 1262 1450 440 395
Citeseer KC 0.47 0.45 0.55 0.64 0.47 0.50 0.24 0.21
$r_{0}$ 1214 1180 1405 1590 1224 1285 405 369
PubMed KC 0.54 0.50 0.62 0.71 0.52 0.66 0.30 0.28
$r_{0}$ 1644 1562 1785 1993 1588 1874 1197 1090
Wiki-CS KC 0.42 0.44 0.40 0.49 0.43 0.45 0.19 0.17
$r_{0}$ 1805 1993 1746 2130 1842 2048 970 904
Amazon-Computers KC 0.39 0.37 0.40 0.45 0.35 0.37 0.12 0.11
$r_{0}$ 1450 1428 1489 1632 1370 1415 874 820
Amazon-Photos KC 0.38 0.38 0.43 0.47 0.39 0.41 0.14 0.12
$r_{0}$ 1872 1884 1990 2145 1895 1921 750 722
Coauthor-CS KC 0.29 0.28 0.32 0.34 0.31 0.32 0.12 0.11
$r_{0}$ 1774 1725 1896 1903 1872 1890 1120 1039
ogbn-arxiv KC 0.12 0.13 0.12 0.14 0.12 0.13 0.05 0.05
$r_{0}$ 1860 1936 1852 1996 1845 1920 1354 1328
5.7 Additional Node Classification Results on More Datasets
Table 8 presents the node classification results under symmetric label noise, asymmetric label noise, and attribute noise on ogbn-arxiv, Wiki-CS, Amazon-Computers, and Amazon-Photos. The table reports the mean accuracy and standard deviation over $10$ independent runs. As shown, both GCL-LRR and GCL-LR-Attention consistently outperform all baseline methods across these benchmark datasets, demonstrating superior robustness to both label and attribute noise.
@@ 5.7 Additional Node Classification Results on More Datasets
Table 8: Performance comparison for node classification on ogbn-arxiv, Wiki-CS, Amazon-Computers, and Amazon-Photos with asymmetric label noise, symmetric label noise, and attribute noise.
Noise Type
@@
GFSA 0.722$\pm$0.006 0.492$\pm$0.012 0.530$\pm$0.012 0.543$\pm$0.010 0.372$\pm$0.012 0.398$\pm$0.011 0.383$\pm$0.011 0.328$\pm$0.016 0.294$\pm$0.015 0.292$\pm$0.012
HONGAT 0.738$\pm$0.005 0.496$\pm$0.010 0.542$\pm$0.011 0.547$\pm$0.009 0.384$\pm$0.013 0.415$\pm$0.012 0.388$\pm$0.012 0.342$\pm$0.015 0.303$\pm$0.014 0.298$\pm$0.011
CRGNN 0.744$\pm$0.004 0.501$\pm$0.009 0.548$\pm$0.010 0.549$\pm$0.008 0.388$\pm$0.011 0.422$\pm$0.010 0.390$\pm$0.010 0.346$\pm$0.014 0.306$\pm$0.013 0.301$\pm$0.010
CGNN 0.732$\pm$0.007 0.494$\pm$0.011 0.538$\pm$0.013 0.545$\pm$0.011 0.378$\pm$0.012 0.408$\pm$0.013 0.385$\pm$0.013 0.335$\pm$0.017 0.300$\pm$0.016 0.296$\pm$0.013
GCL-LRR 0.757$\pm$0.010 0.520$\pm$0.013 0.581$\pm$0.013 0.570$\pm$0.007 0.410$\pm$0.014 0.455$\pm$0.014 0.406$\pm$0.012 0.369$\pm$0.012 0.335$\pm$0.014 0.318$\pm$0.010
GCL-LR-Attention 0.762$\pm$0.010 0.533$\pm$0.013 0.597$\pm$0.013 0.588$\pm$0.007 0.430$\pm$0.014 0.472$\pm$0.014 0.423$\pm$0.012 0.392$\pm$0.012 0.352$\pm$0.014 0.325$\pm$0.010
5.8 Eigen-Projection and Concentration Entropy Analysis on Additional Datasets
Figure 2 presents the eigen-projection visualizations and corresponding signal concentration ratios for Coauthor-CS, Amazon-Computers, Amazon-Photos, and ogbn-arxiv.
5.9 Evaluation on Heterophilic Graphs
In this section, we examine the effectiveness of GCL-LRR for semi-supervised node classification on two commonly used heterophilic graph datasets: Texas and Chameleon [87]. We begin by analyzing the Low Frequency Property (LFP) on these datasets using eigen-projection visualizations and signal concentration ratios, as depicted in Figure 3. The results indicate that the LFP is also present in heterophilic graphs, akin to its presence in homophilic datasets. This analysis is conducted under asymmetric label noise with a corruption level of $60\%$. Using a rank parameter of $0.2\min\left\{N,d\right\}$, the resulting concentration entropy values are $0.762$ for Chameleon and $0.725$ for Texas.
Subsequently, we conduct semi-supervised node classification experiments on Texas and Chameleon following the experimental protocol described in Section 5.2. We employ TEDGCN [88], a widely adopted GNN architecture for heterophilic graphs, as the encoder backbone for both GCL-LRR and GCL-LR-Attention. The classification results are reported in Table 9. As shown, both GCL-LRR and GCL-LR-Attention achieve substantial performance improvements over the base heterophilic GNN, demonstrating robustness to various types of noise in these challenging graph settings.
@@ 5.9 Evaluation on Heterophilic Graphs
Table 9: Performance comparison for node classification on Texas and Chameleon with asymmetric label noise, symmetric label noise, and attribute noise.
Dataset Methods Noise Type
0 40 60 80
- Asymmetric Symmetric Attribute Asymmetric Symmetric Attribute Asymmetric Symmetric Attribute
Texas TEDGCN 0.771$\pm$0.025 0.525$\pm$0.023 0.528$\pm$0.018 0.541$\pm$0.022 0.402$\pm$0.016 0.418$\pm$0.019 0.445$\pm$0.021 0.312$\pm$0.015 0.328$\pm$0.017 0.341$\pm$0.020
GCL-LRR 0.780$\pm$0.013 0.547$\pm$0.019 0.557$\pm$0.016 0.568$\pm$0.017 0.438$\pm$0.015 0.444$\pm$0.017 0.463$\pm$0.018 0.336$\pm$0.012 0.353$\pm$0.014 0.365$\pm$0.016
GCL-LR-Attention 0.785$\pm$0.018 0.556$\pm$0.016 0.563$\pm$0.013 0.576$\pm$0.015 0.451$\pm$0.012 0.452$\pm$0.014 0.472$\pm$0.016 0.338$\pm$0.010 0.367$\pm$0.012 0.372$\pm$0.014
Chameleon TEDGCN 0.569$\pm$0.009 0.382$\pm$0.021 0.401$\pm$0.018 0.425$\pm$0.020 0.298$\pm$0.017 0.315$\pm$0.019 0.328$\pm$0.022 0.225$\pm$0.016 0.241$\pm$0.018 0.254$\pm$0.021
GCL-LRR 0.584$\pm$0.011 0.407$\pm$0.019 0.436$\pm$0.015 0.447$\pm$0.018 0.332$\pm$0.015 0.342$\pm$0.016 0.356$\pm$0.018 0.251$\pm$0.013 0.269$\pm$0.015 0.283$\pm$0.017
GCL-LR-Attention 0.585$\pm$0.008 0.412$\pm$0.016 0.444$\pm$0.013 0.452$\pm$0.014 0.341$\pm$0.011 0.352$\pm$0.013 0.361$\pm$0.015 0.262$\pm$0.010 0.282$\pm$0.012 0.290$\pm$0.014
Figure 3: Eigen-projection (first row) and signal concentration ratio (second row) on Chameleon and Texas. The study in this figure is performed for asymmetric label noise with a noise level of $60\%$. By the rank of $0.2\min\left\{N,d\right\}$, the concentration entropy on Chameleon and Texas are $0.762$ and $0.725$.
5.10 Node Classification Results for GCL Methods with Different Types of Classifiers
Existing GCL approaches, including MERIT [37], SUGRL [38], and SFA [29], typically follow a two-stage procedure: they first train a graph encoder using contrastive objectives such as InfoNCE [37], and subsequently train a linear classifier in a supervised manner on the resulting node representations. In contrast, GCL-LRR integrates a transductive classifier directly atop the contrastively learned representations, enabling label propagation during training. To ensure a fair comparison, we retrain all baseline GCL methods using the same transductive classifier employed in GCL-LRR, as well as an additional two-layer transductive GCN classifier. The results with different types of classifiers are shown in Table 10. It is observed that LRA-GCL still outperforms SOTA GCL methods even when the transductive classifiers are employed.
Table 10: Performance comparison for node classification by inductive linear classifier, transductive two-layer GCN classifier, and transductive classifier used in GCL-LRR. The comparisons are performed on Cora.
Noise Type